# Corresponding author: philipp.vetter(at)uni-ulm.de

### 01 Mar 2024 v1

@@ II.1 The Figures of Merit
Figure 1: Ensemble dynamics and FoM validity measurements. a) Rabi experiment with the N-$V$center ensemble, showing a strong beating. The green areas highlight the time dependence of the Rabi frequency and the black background oscillation visualizes the corresponding shift. b) Linear sweep of the guess pulse amplitude for QPT, c) LGST and $\tilde{G}$, d) RLGST and e) ORBIT. The ordinate shows the absolute value of the FoM as mean value of $20$ measurements. The minimum of the corresponding FoM is indicated by a grey line. f) Comparison of the FoM definitions for the guess (red) and reference pulse (blue), normalized by $\text{FoM}_{guess}$. Each data point shows the mean value of $20$ measurements with the corresponding uncertainty.
@@
for our experiments. Here, $\mathds{1}$ is the identity and $\mathcal{X}$ and $\mathcal{Y}$ are the rotations around the $x$- and $y$- axis, where the rotation angle is given by the index. The gate-set is chosen such that we can generate the full Clifford group [72] needed for RB and ORBIT. We limit our optimizations to $G_{2}=-\mathcal{X}_{\pi/2}$ to analyze how the change of an individual pulse affects the performance of our gate-set Eq. (1) and to ensure that the optimization converges quickly. The initial state is set to $|\rho\rangle\rangle=|0\rangle\rangle$ and the POVM to $|E\rangle\rangle=|0\rangle\rangle$ for all of our experiments. Note that operators and states are given in Hilbert-Schmidt space to benefit from a simplified syntax (see supplementary information [73]).
@@ II.4 Cross-Comparison
Figure 3: Cross-comparison of the closed-loop QOC experiments. a) The x-axis shows the method selected to calculate the FoM during an optimization. We perform five optimization runs per method, represented by five bars, whose color denotes the optimization method. Their height corresponds to the achieved gain for the chosen evaluation method displayed in the y-axis and is is the average of $20$ individual measurements. b) Average gain of the five optimizations per method together with the relative standard deviation. The abscissa corresponds to the method selected for the FoM calculation during the optimization and the ordinate to the chosen evaluation method to calculate the gain. c) The best performing pulse, when evaluated by its optimization method, is re-measured for all other evaluation methods. The error corresponds to the standard deviation of 20 measurements.
Each optimization setting is repeated five times to investigate the stability and reproducibility of the optimization. The cross-evaluation of each experiment is shown in Fig. 3 where a) displays an overview of all obtained gain values after optimization. Each column represents the method used for the FoM calculation during the closed-loop QOC experiment, each row represents the evaluation method and the height of the bars corresponds to the achieved gain, Eq. (2). For the evaluation with randomized benchmarking we use the extracted average error rate [69, 70] instead of a FoM to calculate the corresponding gain. For a better comparison, we also show the average gain over all five optimizations for each method with the corresponding relative standard deviation in Fig. 3 b). In Fig. 3 c) we show the gains of the pulse that performs best upon evaluation with its own optimization method for each of the five runs.
@@ II.8 Optimization with RLGST
Optimizing the pulse shape via RLGST shows a large average gain for all methods. We are able to find an optimized gate-set that reaches gains exceeding a value of $1$ for almost all evaluation methods, as shown by Fig. 3 c). The average value of $0.75$ in Fig. 3 b) and the highest value with $1.13$ when being evaluated with QPT even exceeds the QPT optimization itself. Therefore, not only the individual gate performance greatly increases, but also that of the entire gate-set, which is reflected by the high gain at different time-scales for ORBIT and $\tilde{G}$. Nevertheless, QPT and $\tilde{G}$ show a large variance in the obtained gain, as RLGST considers the performance for longer circuits. Thus, not every optimized pulse shape must necessarily work equally well for a small number of applied pulses.
@@ II.9 Optimization with ORBIT
The optimization with ORBIT shows a large improvement for all evaluation methods. If we take the average over all evaluation methods, ORBIT provides the best performance. Through an optimization with ORBIT, we find gate-sets that perform universally well, regardless of the used evaluation method. For RLGST and $\tilde{G}$ we observe large improvements of the gain, reaching and even surpassing the performance of the reference pulse. The method also shows the highest average gain of all methods when evaluated with RB, which is only just below the reference pulse. Upon evaluation with QPT the optimized pulses on average outperform the ones achieved when being optimized via QPT itself. We observe a large variance between the different optimizations for QPT evaluation, while all other methods exhibit a small variance. Like for the optimization with RLGST, this can be explained by the fact that ORBIT optimizes the full gate-set’s performance while QPT evaluates the single pulse $G_{2}$. Thus, the performance of an individual gate must not be maximized.
@@ III Discussion
Based on our detailed investigation, we can now discuss to what extent the individual methods are suitable for closed-loop optimizations on spin ensemble systems similar to ours. We would like to emphasize that all methods have evidently been able to improve the performance of our gate-set.
@@
One downside of ORBIT is that randomized benchmarking usually requires the use of Clifford gates which are generated by the over-full gate-set, Eq. (1). Additional gates that are to be optimized cannot simply be appended to the gate-set but need to be incorporated into the Clifford gates if possible.
This problem can be circumvented by the optimization via RLGST. Additional gates can simply be added to the gate-set. Optimizations via RLGST achieve similar high gains to ORBIT across all methods, i.e. are able to create a universally well-performing gate-set. Analogous to ORBIT, large circuit lengths are preferred when working with RLGST. At large circuit lengths, RLGST evaluates the system dynamics similar to ORBIT making the method ideal when working with a overfull gate-set as in our case. We show that the method can be used far beyond its intended linear (small error) working regime [67] to define a suitable FoM for closed-loop experiments. Moreover, while not directly investigated in our work, the method also allows to work with an incomplete gate-set [67], which can be of use for experimental systems with limited control. In addition, the method provides estimates of the gate-set and states which can be accessed during the optimization to obtain deeper insights into the system dynamics [75].
In conclusion, through the definition of our FoMs based on RLGST and ORBIT we are able to create a gate-set which performs universally well across all investigated evaluation methods, regardless of their time-scale. We provide a detailed comparison of several gate and gate-set analysis methods and how these can be applied in a closed-loop QOC experiment, tested with a macroscopic ensemble of N-$V$centers. Since the occurring types of error are not exclusive to our system but exist in a wide variety of systems, our work can act as a useful handbook for experimentalists trying to overcome similar problems via QOC.
@@ IV.1 Definitions of the Figures of Merit
Starting with QPT, we measure the probabilities
@@
In contrast to QPT, quantum gate-set tomography [66, 77] takes SPAM errors directly into account by measuring the probabilities of Eq. (5) for all $G_{k}\in\mathcal{G}$:
@@
using the inverse Gram matrix $\widetilde{G}_{0}^{-1}$ to redistribute the errors to all other gates. As the probabilities (7) are gauge invariant, the estimates (8) need to be gauge-corrected. Following the derivations provided in Refs. [66, 65], we then obtain the gauge-transformed LGST estimates $*\hat{G}^{*}_{k},~{}|\hat{\rho}^{*}\rangle\rangle$ and $*|\hat{E}^{*}\rangle\rangle$. Usually, a maximum likelihood estimation (MLE) is used next to find the closest physically correct version of the obtained estimates, i.e., a $\hat{G}^{*}_{k}$ that corresponds to a CPTP map. However, such an MLE is computationally demanding, making it unsuitable for our closed-loop QOC experiments. We therefore use the estimates provided by LGST to evaluate the performance of our gates $G_{k}$ through the difference to their target $T_{k}$:
@@
based on the probability matrices $\tilde{G}_{k}$ of Eq. (7) and their difference to the corresponding target probabilities $\tilde{T}_{k}$. While not providing an estimate, these probabilities are gauge invariant by definition.
@@
RLGST [67] combines the idea of gate-set tomography and randomized circuits to obtain an estimate of the gate-set with little computational overhead within a linear approximation. We start by measuring the probabilities
@@
for $N$ randomly chosen circuits $\mathcal{C}_{i}$ of length $L$ with $i=1...N$. Following the calculations provided in Ref. [67], the probabilities of Eq. (12) are then used to obtain an estimate for the error matrices $e_{j}$, which describe how the measurement deviates from the expected target:
@@ IV.2 Correlation Matrix
with the Pearson correlation coefficient

### 24 Mar 2024 v2

@@ II.1 The Figures of Merit
Figure 1: Ensemble dynamics and FoM validity measurements. a) Rabi experiment with the N-$V$center ensemble, showing a strong beating. The green areas highlight the time dependence of the Rabi frequency and the black background oscillation visualizes the corresponding shift. b) Linear sweep of the guess pulse amplitude for QPT, c) LGST and $\tilde{G}$, d) RLGST and e) ORBIT. The ordinate shows the absolute value of the mean FoM from $20$ measurements and the uncertainty. The minimum of the corresponding FoM is indicated by a grey line. f) Comparison of the FoM definitions for the guess (red) and reference pulse (blue), normalized by $\text{FoM}_{\text{guess}}$. Each data point shows the mean value of $20$ measurements with the corresponding standard deviation.
@@
for our experiments. Here, $\mathds{1}$ is the identity and $\mathcal{X}$ and $\mathcal{Y}$ are the rotations around the $x$- and $y$- axis, where the rotation angle is given by the index. The gate-set is chosen such that we can generate the full Clifford group [72] needed for RB and ORBIT. We limit our optimizations to $G_{2}=-\mathcal{X}_{\pi/2}$ to analyze how the change of an individual pulse affects the performance of our gate-set Eq. (1) and to ensure that the optimization converges quickly. Except for QPT, all methods evaluate the performance of the entire gate-set, leading to the exact same measurements as if we would optimize all gates at once. Thus we do not expect a loss of generality of our results by limiting ourselves to one optimized gate. The initial state is set to $|\rho\rangle\rangle=|0\rangle\rangle$ and the POVM to $|E\rangle\rangle=|0\rangle\rangle$ for all of our experiments. Note that operators and states are given in Hilbert-Schmidt space to benefit from a simplified syntax (see supplementary information [73]).
@@ II.4 Cross-Comparison
Figure 3: Cross-comparison of the closed-loop QOC experiments. a) The x-axis shows the method selected to calculate the FoM during an optimization. We perform five optimization runs per method, represented by five bars, whose color denotes the optimization method. Their height corresponds to the achieved gain for the chosen evaluation method displayed in the y-axis and is is the average of $20$ individual measurements. b) Average gain of the five optimizations per method together with the standard deviation. The abscissa corresponds to the method selected for the FoM calculation during the optimization and the ordinate to the chosen evaluation method to calculate the gain. c) The best performing pulse, when evaluated by its optimization method, is re-measured for all other evaluation methods. The error corresponds to the uncertainty of 20 measurements.
Each optimization setting is repeated five times to investigate the stability and reproducibility of the optimization. The cross-evaluation of each experiment is shown in Fig. 3 where a) displays an overview of all obtained gain values after optimization. Each column represents the method used for the FoM calculation during the closed-loop QOC experiment, each row represents the evaluation method and the height of the bars corresponds to the achieved gain, Eq. (2). For the evaluation with randomized benchmarking we use the extracted average error rate [69, 70] instead of a FoM to calculate the corresponding gain. For a better comparison, we also show the average gain over all five optimizations for each method with the corresponding standard deviation in Fig. 3 b). In Fig. 3 c) we show the gains of the pulse that performs best upon evaluation with its own optimization method for each of the five runs.
@@ II.8 Optimization with RLGST
Optimizing the pulse shape via RLGST shows a large average gain for all methods. We are able to find an optimized gate-set that reaches gains exceeding a value of $1$ for almost all evaluation methods, as shown by Fig. 3 c). The average value of $0.75$ in Fig. 3 b) and the highest value with $1.13$ in Fig. 3 c) when being evaluated with QPT even exceeds the QPT optimization itself. Therefore, not only the individual gate performance greatly increases, but also that of the entire gate-set, which is reflected by the high gain at different time-scales for ORBIT and $\tilde{G}$. Nevertheless, QPT and $\tilde{G}$ show a large variance in the obtained gain relative to the mean value, as RLGST considers the performance for longer circuits. Thus, not every optimized pulse shape must necessarily work equally well for a small number of applied pulses.
@@ II.9 Optimization with ORBIT
The optimization with ORBIT shows a large improvement for all evaluation methods. If we take the average over all evaluation methods, ORBIT provides the best performance. Through an optimization with ORBIT, we find gate-sets that perform universally well, regardless of the used evaluation method. For RLGST and $\tilde{G}$ we observe large improvements of the gain, reaching and even surpassing the performance of the reference pulse. The method also shows the highest average gain of all methods when evaluated with RB, which is only just below the reference pulse. Upon evaluation with QPT the optimized pulses on average outperform the ones achieved when being optimized via QPT itself. We observe a large variance between the different optimizations for QPT evaluation, while all other methods exhibit a small variance relative to their mean value. Like for the optimization with RLGST, this can be explained by the fact that ORBIT optimizes the full gate-set’s performance while QPT evaluates the single pulse $G_{2}$. Thus, the performance of an individual gate must not be maximized.
@@ III Discussion
Based on our detailed investigation, we can now discuss to what extent the individual methods are suitable for closed-loop optimizations with spin ensembles treated as one effective qubit or with qubits in general. We would like to emphasize that all methods have evidently been able to improve the performance of our gate-set.
@@
One downside of ORBIT is that randomized benchmarking usually requires the use of Clifford gates which are generated by the over-full gate-set, Eq. (1). Additional gates that are to be optimized cannot simply be appended to the gate-set but need to be interleaved into the Clifford gates if possible.
This problem can largely be circumvented by the optimization via RLGST. Additional gates can simply be added to the gate-set. Optimizations via RLGST achieve similar high gains to ORBIT across all methods, i.e. are able to create a universally well-performing gate-set. Analogous to ORBIT, large circuit lengths are preferred when working with RLGST. At large circuit lengths, RLGST evaluates the system dynamics similar to ORBIT making the method ideal when working with a overfull gate-set as in our case. We show that the method can be used far beyond its intended linear (small error) working regime [67] to define a suitable FoM for closed-loop experiments. Moreover, while not directly investigated in our work, the method also allows to work with an incomplete gate-set [67], which can be of use for experimental systems with limited control. In addition, the method provides estimates of the gate-set and states which can be accessed during the optimization to obtain deeper insights into the system dynamics [75].
In conclusion, through the definition of our FoMs based on RLGST and ORBIT we are able to create a gate-set which performs universally well across all investigated evaluation methods, regardless of their time-scale. We provide a detailed comparison of several gate and gate-set analysis methods and how these can be applied in a closed-loop QOC experiment, tested with a macroscopic ensemble of N-$V$centers. Since the occurring types of error are not exclusive to our system but exist in a wide variety of systems, our work can act as a useful handbook for experimentalists trying to overcome similar problems via QOC. The methods and the corresponding FoMs that work well for our system are expected to perform similar for other systems with comparable dynamics [59, 76, 77, 78] and will also be applicable to systems with significantly smaller errors [79, 80, 81].
@@ IV.1 Definitions of the Figures of Merit
Starting with QPT, we measure the expectation values
@@
In contrast to QPT, quantum gate-set tomography [66, 83] takes SPAM errors directly into account by measuring the expectation values of Eq. (5) for all $G_{k}\in\mathcal{G}$:
@@
using the inverse Gram matrix $\widetilde{G}_{0}^{-1}$ to redistribute the errors to all other gates. As the expectation values (7) are gauge invariant, the estimates (8) need to be gauge-corrected. Following the derivations provided in Refs. [66, 65], we then obtain the gauge-transformed LGST estimates $*\hat{G}^{*}_{k},~{}|\hat{\rho}^{*}\rangle\rangle$ and $*|\hat{E}^{*}\rangle\rangle$. Usually, a maximum likelihood estimation (MLE) is used next to find the closest physically correct version of the obtained estimates, i.e., a $\hat{G}^{*}_{k}$ that corresponds to a CPTP map. However, such an MLE is computationally demanding, making it unsuitable for our closed-loop QOC experiments. We therefore use the estimates provided by LGST to evaluate the performance of our gates $G_{k}$ through the difference to their target $T_{k}$:
@@
based on the matrices $\tilde{G}_{k}$ of Eq. (7) filled with the measured expectation values and their difference to the corresponding target values $\tilde{T}_{k}$. While not providing an estimate, these expectation values are gauge invariant by definition.
@@
RLGST [67] combines the idea of gate-set tomography and randomized circuits to obtain an estimate of the gate-set with little computational overhead within a linear approximation. We start by measuring the expectation values
@@
for $N$ randomly chosen circuits $\mathcal{C}_{i}$ of length $L$ with $i=1...N$. Following the calculations provided in Ref. [67], the expectation values of Eq. (12) are then used to obtain an estimate for the error matrices $e_{j}$, which describe how the measurement deviates from the expected target:
@@ IV.2 Correlation Matrix
with the correlation coefficient